OpenStudy (anonymous):

|2x-1|=9 pls solve for the solution....provide me with step by step direction

OpenStudy (anonymous):

2x -1 = 9 2x = 10 x = 5 2x - 1 = -9 2x = -8 x = -4

OpenStudy (anonymous):

2x+1>9= to what? pls help me find the answer